### v1

# Covariant effective spacetimes of spherically symmetric electro-vacuum with a cosmological constant

Abstract
An algebraic framework was introduced in our previous works to address the covariance issue in spherically symmetric effective quantum gravity. This paper extends the framework to the electro-vacuum case with a cosmological constant. After analyzing the notion of covariance in the classical theory, we propose an effective Hamiltonian for the electromagnetic field. The effective Hamiltonian together with the effective Hamiltonian constraint of gravity determines an effective dynamical model of gravity coupled to electromagnetic. The resulting model is covariant with respect to both the effective metric and the effective vector potential. By solving the equations of motion derived from the effective Hamiltonian constraint, we obtain several quantum-corrected solutions. Notably, some of these solutions reveal quantum gravity effects manifesting not only in spacetime metrics but also in the electromagnetic field. Finally, the covariance of coupling models with general matter fields is discussed.
I Introduction
As a well-established theory of gravity, general relativity (GR), predicting the formation of black holes (BHs) and describeing the dynamics of the Universe, has successfully passed all experimental and observational tests to date [1, 2]. Nevertheless, the singularities commonly existing according to the singularity theorems [3, 4], due to the geodesic incompleteness of spacetimes, suggest regions where GR breaks down. It is widely expected that a quantum theory of gravity will replace classical GR in these extreme regimes and govern the physics near singularities. As a promising candidate for quantum gravity, loop quantum gravity (LQG) has achieved significant breakthroughs [5, 6, 7, 8, 9, 10, 11, 12]. In particular, both the canonical and covariant formulations of LQG have been individually developed, providing a full understanding of quantum dynamics [13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Moreover, the concepts and techniques developed in full LQG have been successfully applied to symmetry-reduced models, giving rise to the fields of loop quantum cosmology (LQC) and loop quantum black holes (LQBH). Within these frameworks, classical singularities are resolved by some natural scenarios, such as big-bounce cosmology and BH-to-white hole transitions [23, 24, 25, 26, 27, 28, 29, 30, 31].
Despite these achievements, several issues remain, particularly regarding covariance, which poses a challenge for the effective theory of the quantum symmetry-reduced models. In LQC, covariance is less problematic, as the choice of coordinates adapted to the symmetries of homogeneity and isotropy fully resolves the diffeomorphism constraint. However, in LQBH, while the choice of coordinates adapted to spherical symmetry simplifies the constraint structure, it still leaves the radial component of the diffeomorphism constraint unsolved. Thus, it becomes necessary to consider diffeomorphism transformations along a direction that combines both the temporal and radial coordinates. Consequently, the covariance issue in LQBH models becomes non-trivial and requires careful treatments [32, 33, 34]. This difficulty arises because the effective theory of LQBH is formulated within the Hamiltonian framework, which relies on a 3+1 decomposition of spacetime.
There are several approaches to preserving covariance in effective LQBH models. One approach is to express the equations in terms of tensors, similar to classical GR. An example of this approach is the construction of the quantum Oppenheimer-Snyder model [35, 36]. In this model, a spherically symmetric vacuum is glued to a collapsing ball of spherically symmetric, homogeneous, and pressureless dust. The dynamics of the dust ball is governed by the effective Friedmann equation derived from LQC in [25]. It turns out that the exterior spacetime can be uniquely determined using the Israel junction condition [37, 38], under the assumption that the exterior region remains stationary. Since the Israel junction condition is a tensor equation, the resulting quantum-corrected exterior spacetime remains covariant. In other words, one can choose different coordinate systems in the exterior region during the matching process and still obtain the same exterior spacetime metric tensor. Another approach is to introduce a specific matter field to fix the gauge of diffeomorphism [39, 40]. In this approach, different choices of matter fields for gauge fixing can lead to entirely different expressions for the Hamiltonian constraints. Some of these formulations may fail to incorporate key features, such as polymerization, which characterizes LQG. This raises an important question: what kind of matter fields should be selected to ensure that the associated effective Hamiltonian retains these key LQG features? The third approach focuses on constructing an effective Hamiltonian constraint that preserves covariance, as in the classical theory, as done in [41, 42]. More specifically, the effective Hamiltonian constraint is designed to ensure that the gauge transformations of the metric, generated by the constraints, correspond exactly to its spacetime diffeomorphism transformations. In these works, the approach starts with a modified constraint algebra incorporating quantum gravity effects. The covariance equation for the effective Hamiltonian constraint is then systematically derived based on the necessary and sufficient conditions required to maintain covariance. By solving the covariance equation, one can derive various effective Hamiltonian constraints that correspond to different covariant theories. It is important to note that although these works focus on vacuum gravity, the necessary and sufficient conditions for spacetime covariance are applicable to any matter coupling. Therefore, this approach can be extended to study scenarios beyond the vacuum case, such as the dust-coupling model [41, 42]. The aim of this paper is to demonstrate that the framework developed in [41, 42] can also be extended to the electro-vacuum setting with a cosmological constant. This extension will ensure covariance not only for the spacetime metric but also for the electromagnetic (EM) field.
The remainder of this paper is organized as follows. In Sec. II, we first recall the Hamiltonian formulation of gravity coupled to EM field in the presentation of a cosmological constant, then analyze the notion of covariance in the classical case, and present its exact solutions of the classical model. In Sec. III, we turn to the effective model of gravity coupled to EM field. We construct an effective Hamiltonian for EM field, together with the effective Hamiltonian constraint of gravity, ensuring the covariance of effective model. Several exact solutions of effective model are presented. In Sec. IV, the covariance of coupling models with general matter fields is discussed. We show that the resulting spacetime is always covariant for gravity coupled to an arbitrary matter field in effective theory, once the constraint algebra of coupling model has the similar structure as that of the effective vacuum model. In comparison to gravity, the covariance of matter field needs to be concretely treated. Our results are summarized in Sec. V.
II Classical theory
In this section, we will first recall the Hamiltonian formulation of gravity minimally coupled to EM field in the presentation of a cosmological constant. Building on the Hamiltonian framework, we will check the covariance of the model explicitly. Next, we introduce the reduced formulation of the model with the Gauss constraint solved. Using this reduced formulation, we derive the solutions to the equations of motion (EOM) and examine the covariance of the coupling model based on these solutions.
II.1 Hamiltonian framework of classical theory
Let us focus on the spherically symmetric case. The spacetime is based on a 4-D manifold ×$\mathcal{M}\cong\mathbb{R}\times\Sigma$ with $\Sigma\cong{\mathbb{R}}^{+}\times{\mathbb{S}}^{2}$. Here ${\mathbb{R}}^{+}$ denotes the positive real line, and ${\mathbb{S}}^{2}$ is the 2-sphere. Denote the coordinates of the spacetime by $(t,x,\theta,\phi)$. The phase space consists of the canonical variables $(K_{1},E^{1};K_{2},E^{2})$ for gravity and $(\Gamma,P_{\Gamma})$ for EM field [43, 44, 45, 46, 47]. The nontrivial Poisson brackets between these variables read
$\left\{K_{1}(x),E^{1}(y)\right\}$ $=2G\delta(x,y),$ (1)
$\left\{K_{2}(x),E^{2}(y)\right\}$ $=G\delta(x,y),$ (2)
$\left\{\Gamma(x),P_{\Gamma}(y)\right\}$ $=\delta(x,y),$ (3)
with $G$ being the Newtonian constant. The physical meaning of the canonical variables becomes evident from the following facts. The variables $E^{I}$ define the 3-D metric
$q_{ab}{\rm d}x^{a}{\rm d}x^{b}=\frac{(E^{2})^{2}}{E^{1}}({\rm d}x+N^{x}{\rm d}t)^{2}+E^{1}{\rm d}\Omega^{2},$ (4)
where the lowercase latin indices $a,b,\dots=1,2,3$ stand for the spatial indices, and ${\rm d}\Omega^{2}={\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}$. The variable $\Gamma$ describes the spatial component of the vector potential of EM field, namely,
$A_{a}{\rm d}x^{a}=\Gamma{\rm d}x.$ (5)
The dynamics of the coupling model is governed by a set of constraints: the Hamiltonian constraint $H=H^{GR}+H^{EM}$, the diffeomorphism constraint $H_{x}=H_{x}^{GR}+H_{x}^{EM}$, and the Gauss constraint of EM field $\mathcal{G}=G^{EM}$. Here the superscripts “GR” and “EM” stand for gravity (described by GR) and EM field, respectively. Specifically, we have
$H^{GR}$ $:=-\frac{K_{1}K_{2}\sqrt{E^{1}}}{G}-\frac{E^{2}}{2G\sqrt{E^{1}}}-\frac{(K_{2})^{2}E^{2}}{2G\sqrt{E^{1}}}+\frac{(\partial_{x}E^{1})^{2}}{8G\sqrt{E^{1}}E^{2}}$
$\hskip 14.22636pt-\frac{\sqrt{E^{1}}(\partial_{x}E^{1})\partial_{x}E^{2}}{2G(E^{2})^{2}}+\frac{\sqrt{E^{1}}\partial^{2}_{x}E^{1}}{2GE^{2}}+\frac{1}{2G}\sqrt{E^{1}}E^{2}\Lambda,$ (6)
$H^{EM}$ $:=\frac{E^{2}P_{\Gamma}^{2}}{2(E^{1})^{3/2}}+\frac{\sqrt{E^{1}}}{E^{2}}\Gamma G^{EM},$ (7)
$H_{x}^{GR}$ $:=\frac{1}{2G}(2E^{2}\partial_{x}K_{2}-K_{1}\partial_{x}E^{1}),$ (8)
$H_{x}^{EM}$ $:=-\Gamma\partial_{x}P_{\Gamma},$ (9)
$G^{EM}$ $:=-\partial_{x}P_{\Gamma},$ (10)
where $\Lambda$ denotes the cosmological constant. These constraints obey the following algebra:
$\left\{\mathcal{G}[\beta_{1}],\mathcal{G}[\beta_{2}]\right\}$ $=0,$ (11a)
$\left\{\mathcal{G}[\beta],H_{x}[N^{x}]\right\}$ $=-\mathcal{G}[N^{x}\partial_{x}\beta],$ (11b)
$\left\{\mathcal{G}[\beta],H[N]\right\}$ $=-\mathcal{G}\left[N\frac{\sqrt{E^{1}}}{E^{2}}\partial_{x}\beta\right],$ (11c)
$\left\{H_{x}[N^{x}_{1}],H_{x}[N^{x}_{2}]\right\}$ $=H_{x}\left[N^{x}_{1}\partial_{x}N^{x}_{2}-N^{x}_{2}\partial_{x}N^{x}_{1}\right],$ (11d)
$\left\{H_{x}[N^{x}],H[N]\right\}$ $=H\left[N^{x}\partial_{x}N\right],$ (11e)
$\left\{H[N_{1}],H[N_{2}]\right\}$ $=H_{x}\left[S\left(N^{x}_{1}\partial_{x}N^{x}_{2}-N^{x}_{2}\partial_{x}N^{x}_{1}\right)\right],$ (11f)
with the structure function $S=\frac{E^{1}}{(E^{2})^{2}}$. Here we introduced the convention $F[g]\equiv\int{\rm d}xF(x)g(x)$ to denote the smeared constraints. It is worth noting that the present paper adopts the EM Hamiltonian $H^{EM}$ in Eq. (7) that differs from the one provided in [46, 47] by a term proportional to the Gauss constraint. This additional term is introduced to ensure that the Hamiltonian constraint and the diffeomorphism constraint form a subalgebra, as demonstrated in Eqs. (11d)–(11f).
To determine the dynamics in the Hamiltonian framework, we need to choose a lapse function $N$, a shift vector $N^{x}\partial_{x}$, and a Lagrangian multiplier $\Psi$ to define a Hamiltonian
$\mathbb{H}=H[N]+H_{x}[N^{x}]+\mathcal{G}[\Psi].$ (12)
Then, the Hamilton’s equations, as the EOM, read
$\dot{K_{1}}$ :=$:=\left\{K_{1},{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}P_{\Gamma}^{2}+N\frac{G}{\sqrt{E^{1}}E^{2}}\Gamma\mathcal{G},$ (13)
$\dot{K_{2}}$ :=$:=\left\{K_{2},{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}P_{\Gamma}^{2}-N\frac{G\sqrt{E^{1}}}{(E^{2})^{2}}\Gamma\mathcal{G},$ (14)
$\dot{E^{1}}$ :=$:=\left\{E^{1},{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (15)
$\dot{E^{2}}$ :=$:=\left\{E^{2},{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\},$ (16)
$\dot{\Gamma}$ :=$:=\left\{\Gamma,{\mathbb{H}}\right\}=\partial_{x}\Psi+\partial_{x}\left[\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\Gamma\right]+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}P_{\Gamma},$ (17)
$\dot{P_{\Gamma}}$ :=$:=\left\{P_{\Gamma},{\mathbb{H}}\right\}=-\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\mathcal{G},$ (18)
where
${\mathbb{H}}^{GR}:=H^{GR}[N]+H^{GR}_{x}[N^{x}],$ (19)
stands for the gravitational Hamiltonian. Solving the Hamilton’s equations together with the constraint equations $H=0$, $H_{x}=0$ and $\mathcal{G}=0$, we get the solutions $E^{I}(t,x)$, $K_{I}(t,x)$, $\Gamma(t,x)$, and $P_{\Gamma}(t,x)$ as fields on $\mathcal{M}$. Denote the spacetime indicies by the Greek letters $\rho,\sigma,\dots=0,1,2,3$. Then, the 4-D metric $g_{\rho\sigma}$ and 4-D vector potential $A_{\rho}$ are given by [46, 47]
${\rm d}s^{2}$ $=-N^{2}{\rm d}t^{2}+\frac{(E^{2})^{2}}{E^{1}}({\rm d}x+N^{x}{\rm d}t)^{2}+E^{1}{\rm d}\Omega^{2},$ (20)
$A_{\rho}{\rm d}x^{\rho}$ $=\Phi{\rm d}t+\Gamma{\rm d}x,$ (21)
with $\Phi$ relating to $\Psi$ and $\Gamma$ by
$\Phi=\Psi+\left(N\frac{\sqrt{E^{1}}}{E^{2}}+N^{x}\right)\Gamma.$ (22)
II.2 Covariance of classical theory
In this subsection, we will demonstrate the covariance of the classical theory within the Hamiltonian framework algebraically.
It is observed that $H_{x}[N^{x}]$ generates the diffeomorphism transformation along the vector field $N^{x}\partial_{x}$. Therefore, the Hamilton’s equations (13)–(18) can be rewritten by the Lie derivatives as
$\mathcal{L}_{\mathfrak{N}}K_{I}=$$\left\{K_{I},H[N]\right\},$ (23)
$\mathcal{L}_{\mathfrak{N}}E^{I}=$$\left\{E^{I},H[N]\right\},$
$\mathcal{L}_{\mathfrak{N}}\Gamma=$$\left\{\Gamma,H[N]+\mathcal{G}[\Psi]\right\},$
$\mathcal{L}_{\mathfrak{N}}P_{\Gamma}=$$\left\{P_{\Gamma},H[N]+\mathcal{G}[\Psi]\right\},$
where $I=1,2$, $\mathfrak{N}=\partial_{t}-N^{x}\partial_{x}$, and $\mathcal{L}_{\mathfrak{N}}$ denotes the Lie derivative along the vector $\mathfrak{N}$ on $\mathcal{M}$. The fields $K_{1}$, $E^{2}$, $\Gamma$ are scalar densities with weight $1$, and $K_{2}$, $E^{1}$ and $P_{\Gamma}$ are scalars, which is relevant for calculating the Lie derivative. Note that by converting Eqs. (13)–(18) into Eq. (23), we indeed establish the relationship between the evolution of a point in phase space and the corresponding evolution of the fields in spacetime.
Now, let us consider a gauge transformation generated by $H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]$, for arbitrary smeared functions $\alpha$ and $\lambda$, and an arbitrary smeared vector field $\beta^{x}\partial_{x}$. Under the gauge transformation, the phase space variables $W\equiv K_{I},E^{I},\Gamma,P_{\Gamma}$ transform as
$W\rightarrow\tilde{W}\equiv W+\epsilon\delta W+O(\epsilon^{2}),$ (24)
where
$\delta W:=\left\{W,H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\right\}.$ (25)
To show the covariance of the coupling model, we need to demonstrate that the gauge transformations in phase space correspond to the diffeomorphism transformations in spacetime. More precisely, we will prove
$\delta g_{\rho\sigma}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}g_{\rho\sigma},$ (26)
$\delta A_{\rho}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}A_{\rho}+\partial_{\rho}\lambda.$ (27)
Note that in Eq. (27), the extra term $\partial_{\rho}\lambda$, representing an infinitesimal $U(1)$ gauge transformation, arises because $A$ is a gauge field with $U(1)$ gauge symmetry.
Observing that the change $\delta g_{\rho\sigma}$ of $g_{\rho\sigma}$ given by Eq. (20) includes $\delta E^{I}$, $\delta N$ and $\delta N^{x}$, let us first calculate $\delta E^{I}$ by definition (25). Here, as discussed in [41, 42], it is sufficient to consider the case where $\alpha$, $\beta^{x}$ and $\lambda$ are phase space independent. Then, we have
$\delta E^{I}=\left\{E^{I},H[\alpha N]\right\}+\mathcal{L}_{\beta}E^{I},$ (28)
where we used that $\{E^{I}(x),\mathcal{G}(y)\}=0$ and the fact that $H_{x}$ generates the spatial diffeomorphism transformation. Since $H$ does not include the derivatives of $K_{I}$, the first term on the right-hand side of Eq. (28) reduces to
$\left\{E^{I}(x),H[\alpha N]\right\}=\alpha(x)\left\{E^{I},H[N]\right\}=\alpha\mathcal{L}_{\mathfrak{N}}E^{I},$ (29)
where Eq. (23) is used in the last step. Substituting this result into Eq. (28) and noting that $E^{1}$ is a scalar but $E^{2}$ is a scalar density, we finally get
$\delta E^{1}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}E^{1},$ (30)
$\delta E^{2}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}E^{2}-E^{2}\mathfrak{N}^{\rho}\partial_{\rho}\alpha.$
To calculate $\delta N$ and $\delta N^{x}$, we need to consider that the gauge transformation (24) will map a solution $K_{I}(t,x)$, $E^{I}(t,x)$ to a new set of solutions
$\tilde{K}_{I}=$ $K_{I}+\epsilon\left\{K_{I},H[\alpha N]+H_{x}[\beta^{x}]\right\}+O(\epsilon^{2}),$ (31)
$\tilde{E}^{I}=$ $E^{I}+\epsilon\left\{E^{I},H[\alpha N]+H_{x}[\beta^{x}]\right\}+O(\epsilon^{2}).$
The new solutions need to satisfy the Hamilton’s equations with respect to a new lapse function $\tilde{N}=N+\epsilon\delta N$ and a new shift vector $\tilde{N}^{x}=N^{x}+\epsilon\delta N^{x}$. This requirement allows us to derive the expressions for $\delta N$ and $\delta N^{x}$, as shown in [41, 42]. Actually, the works [41, 42], which focused on the vacuum case, derived the transformation law for $N$ and $N^{x}$ based on the two key ingredients: (a) the fact that the diffeomorphism constraint generates spatial diffeomorphism transformations, which results in EOM for $E^{I}$ and $K_{I}$ taking the same forms as Eq. (23); and (b) the structure of constraint algebra, which is the same as Eqs. (11d)–(11f). Clearly, these two ingredients (a) and (b) are preserved for the total constraints in the current model. Thus, the transformation laws for $\delta N$ and $\delta N^{x}$ remain the same as in [41, 42], which read
$\delta N=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}N+N\mathfrak{N}^{\rho}\partial_{\rho}\alpha,$ (32)
$\delta N^{x}=$ $-N^{2}S\partial_{x}\alpha-(\mathcal{L}_{\beta}\mathfrak{N})^{x}.$
Combing Eqs. (30) and (32), we arrive at Eq. (26).
To demonstrate Eq. (27), where $\delta A_{\rho}$ involves $\delta\Gamma$ and $\delta\Psi$, we begin by analyzing $\delta\Gamma$. Similar to getting Eq. (30), we have
$\delta\Gamma=\mathcal{L}_{\alpha\mathfrak{N}+\beta}\Gamma+\left(N\frac{\sqrt{E^{1}}}{E^{2}}\Gamma+\Psi\right)\partial_{x}\alpha-\Gamma\mathfrak{N}^{\rho}\partial_{\rho}\alpha+\partial_{x}\lambda.$ (33)
To calculate $\delta\Psi$, we again consider that the gauge transformation (24) will map a solution $\Gamma(t,x)$, $P_{\Gamma}(t,x)$ to a new set of solutions
$\tilde{\Gamma}=$ $\Gamma+\epsilon\delta\Gamma+O(\epsilon^{2}),$ (34)
$\tilde{P}_{\Gamma}=$ $P_{\Gamma}+\epsilon\delta P_{\Gamma}+O(\epsilon^{2}).$
They need to satisfy the Hamilton’s equations with respect to a new lapse function $\tilde{N}=N+\epsilon\delta N$, a new shift vector $\tilde{N}^{x}=N^{x}+\epsilon\delta N^{x}$, and a new Lagrangian multiplier $\tilde{\Psi}=\Psi+\epsilon\delta\Psi$. Notice that $\delta N$ and $\delta N^{x}$ are given by Eq. (32). Then, this requirement will help us to get $\delta\Psi$, with the detailed derivations shown as follows. Substituting $\tilde{\Gamma}$ and $\tilde{P}_{\Gamma}$ given by Eq. (34) into Eq. (23) and comparing the first order results in $\epsilon$ on both sides, we get
$\mathcal{L}_{\delta\mathfrak{N}}X+\mathcal{L}_{\mathfrak{N}}\delta X=\{X,H[\delta N]+\mathcal{G}[\delta\Psi]\}$ (35)
+$+\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}},$
where $X$ is used to represent either $\Gamma$ or $P_{\Gamma}$. Due to the fact that $\delta\mathfrak{N}=-\delta N^{x}\partial_{x}$, the first term on the left-hand side of Eq. (35) reduces to
$\mathcal{L}_{\delta\mathfrak{N}}X=\left\{X,H_{x}[\delta\mathfrak{N}^{x}]\right\}.$ (36)
Then, we compute the second term on the left-hand side of Eq. (35) and get
$\mathcal{L}_{\mathfrak{N}}\delta X$ (37)
$=$ $\mathcal{L}_{\mathfrak{N}}\Big{\{}X,H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}}$
$=$ $\Big{\{}\left\{X,H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\right\},H[N]+\mathcal{G}[\Psi]\Big{\}}$
+$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
Note that in the last step in Eq. (37), the first term contains the Lie derivatives for the phase-space dependent parts in $\delta X$, and the second term includes the Lie derivative for the phase-space independent variables. Applying the Jacobi identity for Poisson brackets to Eq. (37), we get
$\mathcal{L}_{\mathfrak{N}}\delta X$ (38)
$=$ $\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}}$
+$+\Big{\{}X,\left\{H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda],H[N]+\mathcal{G}[\Psi]\right\}\Big{\}}$
+$+\{X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\}.$
The second term on the right-hand side of Eq. (38) contains Poisson bracket between constraints, and thus it can be evaluated by applying the constraint algebra (11a)–(11f). The result reads
$\mathcal{L}_{\mathfrak{N}}\delta X$ (39)
$=$ $\Big{\{}\left\{X,H[N]+\mathcal{G}[\Psi]\right\},H[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\Big{\}}$
+$+\left\{X,\mathcal{G}\left[N\sqrt{S}(-\Psi\partial_{x}\alpha-\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi\right]\right\}$
$-\Big{\{}X,H_{x}\left[SN^{x}\partial_{x}\alpha\right]\Big{\}}+\Big{\{}X,H[\beta^{x}\partial_{x}N]\Big{\}}$
+$+\Big{\{}X,H[\mathcal{L}_{\mathfrak{N}}(\alpha N)]+H_{x}[\mathcal{L}_{\mathfrak{N}}\beta^{x}]+\mathcal{G}[\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)]\Big{\}}.$
Putting Eqs. (36) and (39) into Eq. (35) yields
$\left\{X,H_{x}\left[\delta\mathfrak{N}^{x}-SN^{2}\partial_{x}\alpha-(\mathcal{L}_{\beta}\mathfrak{N})^{x}\right]\right\}$ (40)
+$+\Big{\{}X,H[\beta^{x}\partial_{x}N+\mathcal{L}_{\mathfrak{N}}(\alpha N)-\delta N]\Big{\}}$
+$+\left\{X,\mathcal{G}\left[\xi\right]\right\}=0,$
where the coefficient $\xi$ is
$\xi=-N\sqrt{S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda)-\delta\Psi.$ (41)
Note that the first two terms in Eq. (40) vanishes due to Eq. (32). Thus, Eq. (40) gives
$\delta\Psi=-N\sqrt{S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda).$ (42)
Thanks to Eqs. (30), (32), (33) and (42), a straightforward computation leads to Eq. (27).
II.3 Reduced formulation
In this subsection, we reduce the degrees of freedom by solving the Gauss constraint, which simplifies the process of obtaining solutions in the next subsection.
Combing the evolution equation (18) with the constraint equation $\mathcal{G}=G^{EM}=0$ leads to that $P_{\Gamma}$ is a constant [46]
$P_{\Gamma}=Q,$ (43)
which is independent on both $t$ and $x$. Moreover, the constraint equation $H_{x}^{EM}=0$ is automatically satisfied due to $H^{EM}_{x}$ being proportional to $\mathcal{G}$. Notice that both the remaining constraints and the right-hand sides of the Hamilton’s equations are independent of $\Gamma$. Hence, we can choose the partial gauge fixing [48]
$\Gamma=0.$ (44)
As a result, Eq. (22) reduces to
$\Psi=\Phi.$ (45)
Equation (17) reads
00 0 $=\partial_{x}\Psi+N\frac{E^{2}}{\left(E^{1}\right)^{3/2}}Q.$ (46)
Now, the pair of $(\Gamma,P_{\Gamma})$ is not dynamical and thus can be dropped out from the canonical variables. Therefore, after solving the Gauss constraint equation $\mathcal{G}=0$ and imposing the partial gauge fixing (44), the phase space of the coupling model consists of $(K_{1},E^{1};K_{2},E^{2})$, and the Hamiltonian constraint $H$ and the diffeomorphism constraint $H_{x}$ reduce to
$\bar{H}$ $=H^{GR}+H^{Q},$ (47)
$\bar{H}_{x}$ $=H_{x}^{GR},$ (48)
where
$H^{Q}\equiv\left.H^{EM}\right|_{P_{\Gamma}=Q}=\frac{E^{2}Q^{2}}{2(E^{1})^{3/2}}.$ (49)
The resulting constraint algebra reads
$\left\{\bar{H}_{x}(N^{x}_{1}),\bar{H}_{x}(N^{x}_{2})\right\}$ =$=\left\{H_{x}^{GR}(N^{x}_{1}),H_{x}^{GR}(N^{x}_{2})\right\}$
$=\bar{H}_{x}(N^{x}_{1}\partial_{x}N^{x}_{2}-N^{x}_{2}\partial_{x}N^{x}_{1}),$ (50a)
$\left\{\bar{H}_{x}(N^{x}],\bar{H}(N)\right\}$ $=\bar{H}(N^{x}\partial_{x}N),$ (50b)
$\left\{\bar{H}(N_{1}),\bar{H}(N_{2})\right\}$ =$=\left\{H^{GR}(N_{1}),H^{GR}(N_{2})\right\}$
$=\bar{H}_{x}\left[S\left(N_{1}\partial_{x}N_{2}-N_{2}\partial_{x}N_{1}\right)\right].$ (50c)
In the first step in Eq. (50c), we notice that $H^{GR}$ does not involve the derivatives of $K_{I}$, while $H^{Q}$ is only a function of $E^{I}$ but not involves the derivatives of $E^{I}$. Thus, $H^{Q}$ does not contribute to the Poisson bracket between the smeared Hamiltonian constraints. It is worth to note that the constraint algebra (50) for the coupling model is precisely the same as the one in the vacuum case. Then, the reduced Hamiltonian
$\bar{\mathbb{H}}$ $:=\int{\rm d}x\left(N\bar{H}+N^{x}\bar{H}_{x}\right),$ (51)
determines the corresponding Hamilton’s equations
$\dot{K_{1}}$ :=$:=\left\{K_{1},\bar{\mathbb{H}}\right\}=\left\{K_{1},{\mathbb{H}}^{GR}\right\}-N\frac{3GE^{2}}{2\left(E^{1}\right)^{5/2}}Q^{2},$ (52)
$\dot{K_{2}}$ :=$:=\left\{K_{2},\bar{\mathbb{H}}\right\}=\left\{K_{2},{\mathbb{H}}^{GR}\right\}+N\frac{G}{2\left(E^{1}\right)^{3/2}}Q^{2},$ (53)
$\dot{E^{1}}$ :=$:=\left\{E^{1},\bar{\mathbb{H}}\right\}=\left\{E^{1},{\mathbb{H}}^{GR}\right\},$ (54)
$\dot{E^{2}}$ :=$:=\left\{E^{2},\bar{\mathbb{H}}\right\}=\left\{E^{2},{\mathbb{H}}^{GR}\right\}.$ (55)
II.4 Classical solutions
In this subsection, by solving Eq. (46), the constraint equations $\bar{H}=0$ and $\bar{H}_{x}=0$ and the Hamilton’s equations (52)-(55) under different gauge fixings, we will obtain the corresponding solutions. Based on the solutions, it is more obvious to understand the covariance from the viewpoint of coordinate transformation.
To this end, let us choose the areal gauge [49, 47, 48]
$E^{1}=x^{2}.$ (56)
Then, the diffeomorphism constraint equation $\bar{H}_{x}=0$ yields
$K_{1}=\frac{E^{2}\partial_{x}K_{2}}{x}.$ (57)
The areal gauge (56) allows us to find the stable solution that satisfies the following stationary condition:
00 0 $=\dot{E^{1}}=\left\{E^{1},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\},$ (58)
00 0 $=\dot{E^{2}}=\left\{E^{2},\bar{H}[N]+\bar{H}_{x}[N^{x}]\right\}.$ (59)
This condition can be used to determine the explicit forms $N$ and $N^{x}$ as
$N$ $=\frac{x}{E^{2}},$ (60)
$N^{x}$ $=-\frac{K_{2}}{E^{2}}x.$ (61)
Inserting Eqs. (56) and (57) into Eq. (47) yields
$\bar{H}=-\frac{E^{2}}{x}\partial_{x}\left\{\frac{x}{2G}\left[1+\frac{GQ^{2}}{x^{2}}-\frac{\Lambda}{3}x^{2}+x^{2}\left(\frac{(K_{2})^{2}}{x^{2}}-\frac{1}{(E^{2})^{2}}\right)\right]\right\}.$ (62)
Hence, the Hamiltonian constraint equation $\bar{H}=0$ implies that the term in the derivative is a constant $M$, i.e.,
$\frac{x}{2G}\left[1+\frac{GQ^{2}}{x^{2}}-\frac{\Lambda}{3}x^{2}+x^{2}\left(\frac{(K_{2})^{2}}{x^{2}}-\frac{1}{(E^{2})^{2}}\right)\right]=M.$ (63)
Furthermore, putting Eqs. (56) and (60) into Eq. (46) yields
00 0 $=\partial_{x}\Psi+\frac{Q}{x^{2}}.$ (64)
It implies
$\Phi=\Psi=\frac{Q}{x},$ (65)
where we used Eq. (45). Thus, the vector potential (21) reduces to
$A_{\rho}{\rm d}x^{\rho}$ $=\frac{Q}{x}{\rm d}t.$ (66)
In what follows, we aim to obtain the explicit solutions in different coordinates. On one hand, we take the Schwarzschild gauge
$N^{x}=0.$ (67)
Inserting Eq. (67) into Eq. (61) yields
$K_{2}$ $=0.$ (68)
Putting Eq. (68) into Eq. (63), we have
$E^{2}$ =±$=\pm\frac{x}{\sqrt{f_{c}(x)}},$ (69)
where
$f_{c}(x)=1-\frac{2GM}{x}+\frac{GQ^{2}}{x^{2}}-\frac{\Lambda}{3}x^{2}.$ (70)
Putting Eq. (69) into Eq. (60) yields
$N^{2}=f_{c}(x).$ (71)
Thus, we obtain the solution in Schwarzschild coordinates with the line element (20) being
${\rm d}s^{2}=-f_{c}(x){\rm d}t^{2}+f_{c}(x)^{-1}{\rm d}x^{2}+x^{2}{\rm d}\Omega^{2}.$ (72)
It is the Reissner-Nordström (RN) solution with a cosmological constant as expected. On the other hand, we take the Painlevé-Gullstrand (PG) gauge
$E^{2}=x.$ (73)
According to Eq. (60), the PG gauge (73) results in
$\ N=1.$ (74)
Inserting Eq. (73) into Eqs. (61) and (63), we get
$N^{x}$ =±$=\pm\sqrt{1-f_{c}(x)}.$ (75)
Then, the solution in PG coordinates is obtained, corresponding to the line element (20) taking
${\rm d}s^{2}$ $=-{\rm d}\tau^{2}+\left[{\rm d}x\pm\sqrt{1-f_{c}(x)}\,{\rm d}\tau\right]^{2}+x^{2}{\rm d}\Omega^{2},$ (76)
where $\tau$ is used to denote the time coordinate in the PG gauge.
These two solutions (72) and (76) are related by the coordinate transformation
${\rm d}t={\rm d}\tau\mp\frac{\sqrt{1-f_{c}(x)}}{f_{c}(x)}\,{\rm d}x,$ (77)
reflecting the fact that the spacetime is covariant. Similarly, the vector potential (66) in these two coordinates can be related by
$A_{\rho}{\rm d}x^{\rho}$ $=\frac{Q}{x}{\rm d}t=\frac{Q}{x}{\rm d}\tau\mp{\rm d}Y(x),$ (78)
where $Y$ is the solution to the ordinary differential equation
$\frac{{\rm d}Y}{{\rm d}x}=\frac{Q}{x}\frac{\sqrt{1-f_{c}(x)}}{f_{c}(x)}.$ (79)
Hence, two vector potentials obtained in different gauges differ only in the gauge term ${\rm d}Y(x)$, implying the covariance of EM field.
III Effective theory
In the above section, we introduce how EM field is coupled to gravity within the Hamiltonian framework of the classical GR. Additionally, the general covariance of the classical model was examined with detailed calculations. Now let us turn to the effective model of EM field coupled to effective quantum gravity proposed by [41, 42], and show its covariance.
III.1 Covariance of effective theory
In [41, 42], although the focus is on the vacuum case, the conclusions can be extended to scenarios with matter coupling. To provide a comprehensive perspective, we first introduce these conclusions in a general setting without excluding matter coupling.
In these works, the diffeomorphism constraint $H_{x}$ which generates the spatial diffeomorphism transformation takes its classical form, while the Hamiltonian constraint is modified into an effective one, denoted by $H_{\rm eff}$, with quantum gravity effects included. They are assumed to satisfy the algebraic relation
$\left\{H_{x}[N^{x}_{1}],H_{x}[N^{x}_{2}]\right\}$ $=H_{x}\left[N^{x}_{1}\partial_{x}N^{x}_{2}-N^{x}_{2}\partial_{x}N^{x}_{1}\right],$ (80a)
$\left\{H_{x}[N^{x}],H_{\rm eff}[N]\right\}$ $=H_{\rm eff}\left[N^{x}\partial_{x}N\right],$ (80b)
$\left\{H_{\rm eff}[N_{1}],H_{\rm eff}[N_{2}]\right\}$ $=H_{x}\left[\mu S\left(N^{x}_{1}\partial_{x}N^{x}_{2}-N^{x}_{2}\partial_{x}N^{x}_{1}\right)\right],$ (80c)
where $\mu$, which depends only on the gravitational degrees of freedom, appears in Eq. (80c) as a result of quantum gravity effects. To match this modified constraint algebra, the classical metric is replaced with an effective metric $g_{\rho\sigma}^{(\mu)}$, given by the line element:
${\rm d}s^{2}=-N^{2}{\rm d}t^{2}+\frac{(E^{2})^{2}}{\mu E^{1}}({\rm d}x+N^{x}{\rm d}t)^{2}+E^{1}{\rm d}\Omega^{2}.$ (81)
With these definitions and assumptions, it is turned out in [41, 42] that the metric $g_{\rho\sigma}^{(\mu)}$ is covariant, meaning that it transforms as
$\delta g_{\rho\sigma}^{(\mu)}$ $=\mathcal{L}_{\alpha\mathfrak{N}+\beta}g_{\rho\sigma}^{(\mu)},$ (82)
where $\delta g_{\rho\sigma}^{(\mu)}$ is the infinitesimal gauge transformation generated by $H_{\rm eff}[\alpha N]+H_{x}[\beta^{x}]$, if and only if the following two conditions are satisfied:
  1. (i)
    $H_{\rm eff}$ is independent of the derivatives of $K_{1}$;
  2. (ii)
    For all phase-space-independent functions $\alpha(x)$ and $N(x)$, the following condition holds:
    $\{\mu(x)S(x),H_{\rm eff}[\alpha N]\}=\alpha(x)\{\mu(x)S(x),H_{\rm eff}[N]\}.$ (83)
In the vacuum case, $H_{\rm eff}$ is just the effective Hamiltonian constraint of gravity, denoted by $H_{\rm eff}^{GR}$. With restricting the number of terms for consideration, it is derived from the above two conditions (i) and (ii) that $H_{\rm eff}^{GR}$ must be given by the following equations [41, 42]
$H_{\rm eff}^{GR}=$ $-2E^{2}\left[\partial_{s_{1}}M_{\rm eff}+\frac{\partial_{s_{2}}M_{\rm eff}}{2}s_{3}+\frac{\partial_{s_{4}}M_{\rm eff}}{s_{4}}s_{5}+\mathcal{R}(s_{1},M_{\rm eff})\right],$ (84)
where $\mathcal{R}$ is an arbitrary function of $s_{1}$ and $M_{\rm eff}(s_{1},s_{2},s_{4})$, and $M_{\rm eff}(s_{1},s_{2},s_{4})$ satisfies the following equations evolving a function $\mu(s_{1},s_{2},s_{4})$:
$\frac{\mu s_{1}s_{4}}{4G^{2}}=\left(\partial_{s_{2}}M_{\rm eff}\right)\partial_{s_{2}}\partial_{s_{4}}M_{\rm eff}-\left(\partial_{s_{4}}M_{\rm eff}\right)\partial_{s_{2}}^{2}M_{\rm eff},$ (85a)
$\left(\partial_{s_{2}}\mu\right)\partial_{s_{4}}M_{\rm eff}-\left(\partial_{s_{2}}M_{\rm eff}\right)\partial_{s_{4}}\mu=0.$ (85b)
Here $s_{a}$ are spacetime scalars defined by
$s_{1}=E^{1},\qquad\ s_{2}=K_{2},\qquad s_{3}=\frac{K_{1}}{E^{2}},$ (86)
$s_{4}=\frac{\partial_{x}E^{1}}{E^{2}},\qquad s_{5}=\frac{1}{E^{2}}\partial_{x}\left(\frac{\partial_{x}E^{1}}{E^{2}}\right).$ (87)
It is easy to show that the classical Hamiltonian constraint (6) can be recovered if we set the free function $\mathcal{R}$ to be
$\mathcal{R}_{\rm cl}$ $=-\frac{\Lambda}{4G}\sqrt{s_{1}},$ (88)
and take
$M_{\rm eff}\equiv M_{\rm cl}$ $=\frac{\sqrt{s_{1}}}{2G}\left[1+(s_{2})^{2}-\frac{(s_{4})^{2}}{4}\right],$ (89)
which is a solution to Eq. (85) with $\mu=1$. In this paper, to incorporate the quantum gravity effects, we will take
$\mathcal{R}$ $=-\mathcal{X}(\mu)\frac{\Lambda}{4G}\sqrt{s_{1}},$ (90)
where $\mathcal{X}$ denotes an arbitrary function of $\mu$.
For the EM field coupling, the covariance of $g_{\rho\sigma}^{(\mu)}$ must be preserved. This requires modifying the classical EM Hamiltonian into an effective form, $H_{\rm eff}^{EM}$, to ensure that the constraint algebra (80) and the covariance conditions (i) and (ii) remain valid for
$H_{\rm eff}=H_{\rm eff}^{GR}+H_{\rm eff}^{EM}.$ (91)
To achieve this, we introduce
$H^{EM}_{\rm eff}=\mathcal{Y}(\mu)\frac{E^{2}}{2(E^{1})^{3/2}}P_{\Gamma}^{2}-\frac{\sqrt{\mu E^{1}}}{E^{2}}\Gamma\partial_{x}P_{\Gamma},$ (92)
with an arbitrary function $\mathcal{Y}(\mu)$. Using Eq. (85), it can be verified through straightforward calculations that the algebra generated by the total constraints reads
$\{\mathcal{G}[\beta_{1}],\mathcal{G}[\beta_{2}]\}$ $=0,$ (93a)
$\{\mathcal{G}[\beta],H_{x}[N]\}$ $=-\mathcal{G}[N^{x}\partial_{x}\beta],$ (93b)
$\{\mathcal{G}[\beta],H_{\rm eff}[N]\}$ $=-\mathcal{G}\left[N\frac{\sqrt{\mu E^{1}}}{E^{2}}\partial_{x}\beta\right],$ (93c)
$\{H_{x}[N^{x}_{1}],H_{x}[N^{x}_{2}]\}$ $=H_{x}[N^{x}_{1}\partial_{x}N^{x}_{2}-N_{2}^{x}\partial_{x}N_{1}^{x}],$ (93d)
$\{H_{x}[N^{x},H_{\rm eff}[N]]\}$ $=H_{\rm eff}[N^{x}\partial_{x}N],$ (93e)
$\{H_{\rm eff}[N_{1}],H_{\rm eff}[N_{2}]\}$ $=H_{x}[\mu S(N_{1}\partial_{x}N_{2}-N_{2}\partial_{x}N_{1})].$ (93f)
Hence, the constraints $H_{x}$ and $H_{\rm eff}$ form a subalgebra as proposed by Eq. (80). Now let us show that the covariance conditions (i) and (ii) are satisfied for $H_{\rm eff}$. At first, the derivatives of $K_{1}$ are not contained in $H_{\rm eff}$ as that for $H_{\rm eff}^{EM}$, ensuring that the condition (i) is satisfied. Moreover, a direct calculation shows that
$\left\{\mu,H_{\rm eff}^{EM}[\alpha N]\right\}=\alpha\left\{\mu,H_{\rm eff}^{EM}[N]\right\}.$ (94)
Since the derivatives of $K_{I}$ are not included in $H_{\rm eff}^{EM}$, Eq. (94) implies
$\left\{\mu S,H_{\rm eff}^{EM}[\alpha N]\right\}=\alpha\left\{\mu S,H_{\rm eff}^{EM}[N]\right\}.$ (95)
As this relation must hold for $H_{\rm eff}^{GR}$ due to the covariance of the vacuum model, the condition (ii) is thus satisfied.
Now let us turn to the covariance of EM field in effective case. Due to the factor $\mu$ in the constraint algebra (93), it is easy to show following the procedure given in Sec. II.2 that the vector potential $A_{\rho}$ is no longer covariant, i.e., Eq. (27) does not hold any more. To fix the problem, as done for the metric, we will introduce the effective vector potential, denoted by $A_{\rho}^{(\mu)}$, to restore covariance of EM field. In other words, we will define $A_{\rho}^{(\mu)}$ such that
$\delta A_{\rho}^{(\mu)}=\mathcal{L}_{\alpha\mathfrak{N}+\beta}A_{\rho}^{(\mu)}+\partial_{\rho}\lambda.$ (96)
To get the correct expression of $A_{\rho}^{(\mu)}$, we need to repeat the derivation shown in Sec. II.2, but with the new constraint algebra (93). In this case, Eqs. (32), (33) and (42) become
$\delta N=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}N+N\mathfrak{N}^{\rho}\partial_{\rho}\alpha,$ (97)
$\delta N^{x}=$ $-N^{2}\mu S\partial_{x}\alpha-(\mathcal{L}_{\beta}\mathfrak{N})^{x},$
$\delta\Gamma=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}\Gamma+\left(N\frac{\sqrt{\mu E^{1}}}{E^{2}}\Gamma+\Psi\right)\partial_{x}\alpha-\Gamma\mathfrak{N}^{\rho}\partial_{\rho}\alpha+\partial_{x}\lambda,$
$\delta\Psi=$ $-N\sqrt{\mu S}(\Psi\partial_{x}\alpha+\partial_{x}\lambda)+\beta^{x}\partial_{x}\Psi+\mathcal{L}_{\mathfrak{N}}(\alpha\Psi+\lambda).$
Moreover, $\delta E^{I}$ in Eq. (30) remains unchanged. Since Eq. (94) also holds for $H_{\rm eff}$, we have
$\delta\mu=$$\{\mu,H_{\rm eff}[\alpha N]+H_{x}[\beta^{x}]+\mathcal{G}[\alpha\Psi+\lambda]\}$ (98)
$=$ $\alpha\mathcal{L}_{\mathfrak{N}}\mu+\mathcal{L}_{\beta}\mu$
$=$ $\mathcal{L}_{\alpha\mathfrak{N}+\beta}\mu,$
where we have used the EOM (23) with $H$ replaced by $H_{\rm eff}$ in the second step, and the fact that $\mu$ is a spacetime scalar in the third step. Combining all these results, we obtain an explicit form of $A_{\rho}^{(\mu)}$ as
$A_{\rho}^{(\mu)}{\rm d}x^{\rho}=\left[\Psi+\left(N\frac{\sqrt{\mu E^{1}}}{E^{2}}+N^{x}\right)\Gamma\right]{\rm d}t+\Gamma{\rm d}x,$ (99)
such that Eq. (96) is satisfied.
III.2 Effective dynamics
Let us follow the procedure introduced in the Secs. II.3 and II.4 to solve the dynamics of the effective model, determined by $H_{\rm eff}^{GR}$ in Eq. (84) with $\mathcal{R}$ given by (90) and $H_{\rm eff}^{EM}$ in Eq. (92). In this case of $\mathcal{R}\neq 0$, $M_{\rm eff}$ is no longer a Dirac observable, while it is in the case of $\mathcal{R}=0$ for the vacuum model without a cosmological constant considered in [41, 42].
Similar as in Sec. II.3, the Gauss constraint $\mathcal{G}=0$ implies
$P_{\Gamma}=Q.$ (100)
Then, choosing the condition $\Gamma=0$ to fix the gauge transformation generated by the Gauss constraint, we get the EOM for $\Psi$ similar to Eq. (46) as
$0=\partial_{x}\Psi+\mathcal{Y}(\mu)\frac{NE^{2}}{(E^{1})^{3/2}}Q.$ (101)
With this result, the effective Hamiltonian constraint and the diffeomorphism constraint become
$\bar{H}_{\rm eff}=$ $H_{\rm eff}^{GR}+H_{\rm eff}^{Q},$ (102)
$\bar{H}_{x}=$ $H_{x}^{GR},$
where we introduced
$H_{\rm eff}^{Q}=H_{\rm eff}^{EM}\Big{|}_{P_{\Gamma}=Q}=\mathcal{Y}(\mu)\frac{E^{2}Q^{2}}{2(E^{1})^{3/2}}.$ (103)
The reduced diffeomorphism constraint $\bar{H}_{x}$ is solved by imposing the areal gauge condition (56), leading once again to the solution given in Eq. (57). Then, as in Sec. II.4, we are allowed to find the stable solution satisfying the stationary condition
00 0 $=\dot{E^{1}}=\left\{E^{1},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\},$ (104)
00 0 $=\dot{E^{2}}=\left\{E^{2},\bar{H}_{\rm eff}[N]+\bar{H}_{x}[N^{x}]\right\}.$
Under the areal gauge (56) and (57), substituting Eq. (102) into Eq. (104), we have
$0=$ $GNM_{{\rm eff},s_{2}}+xN^{x},$ (105)
$0=$ $\partial_{x}\left(N^{x}E^{2}\right)+\frac{GNE^{2}}{x}\partial_{x}M_{{\rm eff},s_{2}}$
$-\frac{NE^{2}M_{{\rm eff},s_{2}}}{2x^{3}}\left(x^{4}\Lambda\mathcal{X}_{M}+GQ^{2}\mathcal{Y}_{M}\right),$
with the abbreviations
$M_{{\rm eff},s_{2}}\equiv\partial_{s_{2}}M_{\rm eff},\quad\mathcal{X}_{M}\equiv\frac{{\rm d}\mathcal{X}}{{\rm d}\mu}\frac{\partial\mu}{\partial M_{\rm eff}},\quad\mathcal{Y}_{M}\equiv\frac{{\rm d}\mathcal{Y}}{{\rm d}\mu}\frac{\partial\mu}{\partial M_{\rm eff}}.$ (106)
Here, it is used that $\mu$ is a function of $s_{1}$ and $M_{\rm eff}$ as a consequence of (85b). Equation (105) can be further simplified as
$N^{x}$ $=-\frac{GM_{{\rm eff},s_{2}}}{x}N,$ (107a)
$N$ $=\frac{x}{E^{2}}\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{2Gx^{2}}{\rm d}x\right].$ (107b)
It should be emphasized that the functions $M_{{\rm eff},s_{2}}$, $\mathcal{X}_{M}$ and $\mathcal{Y}_{M}$ in Eqs. (105) and (107) need to be evaluated under the areal gauge (56) and (57).
Now let us consider the effective Hamiltonian constraint. In the areal gauge (56) and (57), $\bar{H}_{\rm eff}=0$ can be rewritten as
$\bar{H}_{\rm eff}=-\frac{E^{2}}{x}\left[\partial_{x}M_{\rm eff}-\frac{x^{4}\Lambda\mathcal{X}(\mu)+GQ^{2}\mathcal{Y}(\mu)}{2Gx^{2}}\right]=0.$ (108)
Under the gauge condition (56) and (57), $M_{\rm eff}$, $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$ become functions of $E^{2}$ and $K_{2}$. In other words, Eq. (108) establishes a relation between $K_{2}$ and $E^{2}$. To solve this equation, a new gauge, such as the Schwarzschild gauge (67) or the PG gauge (73), is needed. Once such a gauge is imposed, Eq. (108) can be solved, and $N$ and $N^{x}$ can be determined using Eq. (107). With $\mu$ obtained from these solutions, the effective metric (81) then can be constructed.
To get the solution for the effective vector potential $A_{\rho}^{(\mu)}$, we need to solve Eq. (101). Under the areal gauge, using Eq. (107b), Eq. (101) reduces to
$\partial_{x}\Psi+\frac{Q}{x^{2}}\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{2Gx^{2}}{\rm d}x\right]\mathcal{Y}(\mu)=0.$ (109)
Solving this equation to get $\Psi_{\rm eff}$, we finally have
$A_{\rho}^{(\mu)}{\rm d}x^{\rho}=\Psi_{\rm eff}(x){\rm d}t,$ (110)
where we have chosen $\Gamma=0$.
In principle, the results depend on the choices of $M_{\rm eff}$ and $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$. In [41, 42], three solutions $M_{\rm eff}$ to Eq. (85) are proposed, with incorporating polymerization for the classical expression of $M_{\rm eff}$. For convenience, we denote these three solutions by subscripts I, II, and III, respectively. Namely, these three solutions will be denoted by $M_{\rm eff,I}$, $M_{\rm eff,II}$, and $M_{\rm eff,III}$. In what follows, let us consider the resulting effective metrics and vector potentials.
III.3 Solutions to effective dynamics from $M_{\rm eff,I}$
Specifically, the expression for $M_{\rm eff,I}$ is [41, 42]
$M_{\rm eff,I}=$ $\frac{\sqrt{s_{1}}}{2G}\left[1+\frac{s_{1}}{\zeta^{2}}\sin^{2}\left(\frac{\zeta s_{2}}{\sqrt{s_{1}}}\right)-\frac{(s_{4})^{2}}{4}\exp\left(\frac{\mathrm{i}2\zeta s_{2}}{\sqrt{s_{1}}}\right)\right]$
$=$ $\frac{\sqrt{E^{1}}}{2G}+\frac{(E^{1})^{3/2}\sin^{2}\left(\frac{\zeta K_{2}}{\sqrt{E^{1}}}\right)}{2G\zeta^{2}}$
$-\frac{\sqrt{E^{1}}}{8G}\left(\frac{\partial_{x}E^{1}}{E^{2}}\right)^{2}\exp\left(\frac{\mathrm{i}2\zeta K_{2}}{\sqrt{E^{1}}}\right).$ (111)
Here $\zeta$ denotes a quantum parameter. Putting Eq. (III.3) into the covariance equations (85), we obtain the corresponding $\mu$ as
$\mu_{\rm I}=1.$ (112)
As a consequence, $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ are two constants. Without loss of generality, we set $\mathcal{X}(\mu_{\rm I})=\mathcal{Y}(\mu_{\rm I})=1$, since any other non-vanishing values of $\mathcal{X}(\mu_{\rm I})$ and $\mathcal{Y}(\mu_{\rm I})$ would simply correspond to redefinitions of $\Lambda$ and $Q$.
Inserting $\mathcal{X}(\mu)=\mathcal{Y}(\mu_{\rm I})=1$ into Eq. (108), we get
$M_{\rm eff,I}=M-\frac{Q^{2}}{2x}+\frac{\Lambda x^{3}}{6G},$ (113)
with $M$ being the integration constant. Due to the fact that $\mathcal{X}(\mu_{\rm I})=\mathcal{Y}(\mu_{\rm I})=1$ leads to $\mathcal{X}_{M}=\mathcal{Y}_{M}=0$, Eq. (107b) gives
$N=\frac{x}{E^{2}}.$ (114)
Then, Eq. (109) can be solved easily to get
$\Psi_{\rm eff,I}(x)=\frac{Q}{x}.$ (115)
Therefore, the effective vector potential (110) reads
$A_{\rho}^{(\mu_{\rm I})}{\rm d}x^{\rho}=\frac{Q}{x}{\rm d}t.$ (116)
Now, we need to impose an additional gauge fixing condition. Let us choose the Schwarzschild gauge (67), which together with Eq. (107a) implies
$-x^{2}\left[\frac{{\rm i}\zeta e^{\frac{{\rm i}2\zeta K_{2}}{x}}}{(E^{2})^{2}}-\frac{\sin\left(\frac{2\zeta K_{2}}{x}\right)}{2\zeta}\right]=0.$ (117)
Combing Eq. (117) with Eq. (III.3), one can eliminate $K_{2}$ to get
$E^{2}$ =±$=\pm\frac{x}{\sqrt{\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\left[1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\right]}}.$ (118)
Plugging Eq. (118) into Eq. (114), we obtain
$N^{2}$ $=\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\left[1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\right].$ (119)
Hence, in Schwarzschild coordinates, the effective line element (81) takes the form
${\rm d}s^{2}_{\rm eff,I}$ $=-f_{\rm eff,I}(x){\rm d}t^{2}+f_{\rm eff,I}(x)^{-1}{\rm d}x^{2}+x^{2}{\rm d}\Omega^{2},$ (120)
where
$f_{\rm eff,I}(x)$ $=\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\left[1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,I}}{x}\right)\right]$
$=f_{c}(x)\left[1+\frac{\zeta^{2}}{x^{2}}f_{c}(x)\right].$ (121)
The solution (120) reduces to the one obtained in [50] as $Q\rightarrow 0$, and to the vacuum one in [41, 42] as $Q=\Lambda\rightarrow 0$. We can also choose the PG gauge (73) to get the metric. In this gauge, the lapse function becomes to $N=1$, according to Eq. (114). Combining Eq. (107a) with Eq. (113) yields
$N^{x}=\pm\sqrt{1-f_{\rm eff,I}(x)}.$ (122)
Thus, in PG coordinates, the effective line element (81) is
${\rm d}s^{2}_{\rm eff,I}$ $=-{\rm d}\tau^{2}+\left[{\rm d}x\pm\sqrt{1-f_{\rm eff,I}(x)}\,{\rm d}\tau\right]^{2}+x^{2}{\rm d}\Omega^{2}.$ (123)
It is easy to see that these two solutions (120) and (123) are related by the coordinate transformation
${\rm d}t={\rm d}\tau\mp\frac{\sqrt{1-f_{\rm eff,I}(x)}}{f_{\rm eff,I}(x)}\,{\rm d}x.$ (124)
Then, the vector potential (110) in these two coordinates can be related by
$A_{\rho}^{(\mu_{\rm I})}{\rm d}x^{\rho}$ $=\frac{Q}{x}{\rm d}t=\frac{Q}{x}{\rm d}\tau\mp{\rm d}Y_{\rm eff,I}(x),$ (125)
where $Y_{\rm eff,I}$ is the solution to the ordinary differential equation
$\frac{{\rm d}Y_{\rm eff,I}}{{\rm d}x}=\frac{Q}{x}\frac{\sqrt{1-f_{\rm eff,I}(x)}}{f_{\rm eff,I}(x)}.$ (126)
Hence, these two vector potentials obtained in these two different gauges are the same up to the gauge term ${\rm d}Y_{\rm eff,I}(x)$, reflecting the covariance of EM field.
III.4 Solutions to effective dynamics from $M_{\rm eff,II}$
The second solution $M_{\rm eff,II}$ to Eq. (85) reads [41, 42]
$M_{\rm eff,II}=$ $\frac{\sqrt{s_{1}}}{2G}\left[1+\frac{s_{1}}{\zeta^{2}}\sin^{2}\left(\frac{\zeta s_{2}}{\sqrt{s_{1}}}\right)-\frac{(s_{4})^{2}}{4}\cos^{2}\left(\frac{\zeta s_{2}}{\sqrt{s_{1}}}\right)\right]$
$=$ $\frac{\sqrt{E^{1}}}{2G}+\frac{(E^{1})^{3/2}\sin^{2}\left(\frac{\zeta K_{2}}{\sqrt{E^{1}}}\right)}{2G\zeta^{2}}$
$-\frac{\sqrt{E^{1}}}{8G}\left(\frac{\partial_{x}E^{1}}{E^{2}}\right)^{2}\cos^{2}\left(\frac{\zeta K_{2}}{\sqrt{E^{1}}}\right).$ (127)
The corresponding $\mu$ reads
$\mu_{\rm II}$ $=1+\frac{\zeta^{2}}{E^{1}}\left(1-\frac{2GM_{\rm eff,II}}{\sqrt{E^{1}}}\right)$
$=1+\frac{\zeta^{2}}{x^{2}}\left(1-\frac{2GM_{\rm eff,II}}{x}\right).$ (128)
To determine resulting solution, one needs to make a certain gauge fixing. In the Schwarzschild gauge (67), Eq. (107a) leads to
$x^{2}\frac{\zeta^{2}+(E^{2})^{2}}{2\zeta(E^{2})^{2}}\sin\left(\frac{2\zeta K_{2}}{x}\right)=0.$ (129)
Combing Eqs. (129) and (III.4), we have
$E^{2}$ =±$=\pm\frac{x}{\sqrt{1-\frac{2GM_{\rm eff,II}}{x}}}$
±$\equiv\pm\frac{x}{\sqrt{f_{\rm eff,II}(x)}}.$ (130)
Using Eq. (107b), we get
$N^{2}$ $=\left(\frac{x}{E^{2}}\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{2Gx^{2}}{\rm d}x\right]\right)^{2}$
$=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]f_{\rm eff,II}(x)$
$\equiv e^{g_{\rm eff,II}(x)}f_{\rm eff,II}(x).$ (131)
Then, the effective line element (81) reads
${\rm d}s^{2}$ $=-e^{g_{\rm eff,II}(x)}f_{\rm eff,II}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm II}(x)f_{\rm eff,II}(x)}+x^{2}{\rm d}\Omega^{2}.$ (132)
Similarly, in the PG gauge (73), combing Eqs. (107a) and (III.4), we have
$N^{x}$ =±$=\pm N\sqrt{\mu_{\rm II}\frac{2GM_{\rm eff,II}}{x}}$
=±$=\pm N\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}.$ (133)
Using Eq. (107b), we have
$N^{2}$ $=\left(\frac{x}{E^{2}}\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{2Gx^{2}}{\rm d}x\right]\right)^{2}$
$=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]$
$=e^{g_{\rm eff,II}(x)}.$ (134)
Thus, the effective line element (81) reduces to
${\rm d}s^{2}=$ $-e^{g_{\rm eff,II}(x)}{\rm d}\tau^{2}$
$+\frac{1}{\mu_{\rm II}(x)}\left({\rm d}x\pm\sqrt{e^{g_{\rm eff,II}(x)}\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}{\rm d}\tau\right)^{2}$
$+x^{2}{\rm d}\Omega^{2}.$ (135)
These two solutions (132) and (III.4) are related by the coordinate transformation
${\rm d}t={\rm d}\tau\mp\frac{\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}}{e^{\frac{1}{2}g_{\rm eff,II}(x)}\mu_{\rm II}(x)f_{\rm eff,II}(x)}\,{\rm d}x,$ (136)
reflecting the fact that the effective spacetime is covariant.
On the other hand, Eq. (109) can be written as.
$\partial_{x}\Psi_{\rm eff,II}+\frac{Q}{x^{2}}e^{\frac{1}{2}g_{\rm eff,II}(x)}\mathcal{Y}(\mu_{\rm II})=0.$ (137)
Hence, the solution to Eq. (137) is can be expressed as
$\Psi_{\rm eff,II}(x)=-Q\int\frac{e^{\frac{1}{2}g_{\rm eff,II}(x)}}{x^{2}}\mathcal{Y}(\mu_{\rm II}){\rm d}x,$ (138)
where the quantum corrections are incorporated into the expression. Note that the solution $\Psi_{\rm eff,II}(x)$ to Eq. (137) depends only on $x$. Thus
$A_{\rho}^{(\mu_{\rm II})}{\rm d}x^{\rho}$ $=\Psi_{\rm eff,II}(x){\rm d}t$
$=\Psi_{\rm eff,II}(x)\left[{\rm d}\tau\mp\frac{\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}}{e^{\frac{1}{2}g_{\rm eff,II}(x)}\mu_{\rm II}(x)f_{\rm eff,II}(x)}\,{\rm d}x\right]$
$=\Psi_{\rm eff,II}(x){\rm d}\tau\mp{\rm d}Y_{\rm eff,II}(x),$ (139)
where
$\frac{{\rm d}Y_{\rm eff,II}(x)}{{\rm d}x}=\Psi_{\rm eff,II}(x)\frac{\sqrt{\mu_{\rm II}(x)\left[1-f_{\rm eff,II}(x)\right]}}{e^{\frac{1}{2}g_{\rm eff,II}(x)}\mu_{\rm II}(x)f_{\rm eff,II}(x)},$ (140)
indicating the covariance of EM field.
In the following, we will calculate the values $M_{\rm eff,II}$, $g_{\rm eff,II}(x)$ and $\Psi_{\rm eff,II}(x)$ for different choices of $\mathcal{X}(\mu_{\rm II})$ and $\mathcal{Y}(\mu_{\rm II})$.
III.4.1 Case of $\mathcal{X}(\mu_{\rm II})=\mathcal{Y}(\mu_{\rm II})=1$
The simplest case is to choose
$\mathcal{X}(\mu_{\rm II})=1,\quad\mathcal{Y}(\mu_{\rm II})=1.$ (141)
In this case, we have
$\mathcal{X}_{M}=0,\quad\mathcal{Y}_{M}=0.$ (142)
Thus, the factor $g_{\rm eff,II}$ takes
$g_{\rm eff,II}^{(1)}=0.$ (143)
Equation (108) can be solved easily with the solution
$M_{\rm eff,II}^{(1)}\equiv M_{\rm eff,II}=M-\frac{Q^{2}}{2x}+\frac{\Lambda x^{3}}{6G},$ (144)
where $M$ denotes the integration constant. Then, Eq. (138) reduces to
$\Psi_{\rm eff,II}^{(1)}(x)=\frac{Q}{x}.$ (145)
III.4.2 Case of $\mathcal{X}(\mu_{\rm II})=1$ and $\mathcal{Y}(\mu_{\rm II})=\mu_{\rm II}$
It will be interesting to consider a case of $\mathcal{X}(\mu_{\rm II})=1$ and $\mathcal{Y}(\mu_{\rm II})\neq 1$. An analytically solvable case is to take
$\mathcal{X}(\mu_{\rm II})=1,\quad\mathcal{Y}(\mu_{\rm II})=\mu_{\rm II}.$ (146)
In this case, we get
$\mathcal{X}_{M}=0,\quad\mathcal{Y}_{M}=-\frac{2\zeta^{2}G}{x^{3}}.$ (147)
Then, the factor $g_{\rm eff,II}$ reads
$g_{\rm eff,II}^{(2)}(x)=-\frac{\zeta^{2}GQ^{2}}{4x^{4}}.$ (148)
Putting Eq. (146) into Eq. (108), we get
$M_{\rm eff,II}^{(2)}=$ $\frac{\Lambda x^{3}}{6G}+e^{\frac{\zeta^{2}GQ^{2}}{4x^{4}}}\left[\frac{\zeta^{2}Q^{2}}{8x^{3}}E_{\frac{1}{4}}\left(\frac{\zeta^{2}GQ^{2}}{4x^{4}}\right)\right.$
$+\frac{Q^{2}}{8x}\left(1-\frac{\zeta^{2}\Lambda}{3}\right)E_{\frac{3}{4}}\left(\frac{\zeta^{2}GQ^{2}}{4x^{4}}\right)$
$\left.+M-\frac{Q^{3/2}\mathbf{\Gamma}\left(\frac{1}{4}\right)}{4\sqrt{2\zeta}\sqrt[4]{G}}\right],$ (149)
where $M$ is the integration constant, $E_{\alpha}(x)=\int_{1}^{\infty}e^{-xt}/t^{\alpha}{\rm d}t$ is the exponential integral function, and $\mathbf{\Gamma}$ denotes the Gamma function. The constant $\frac{Q^{3/2}\Gamma\left(\frac{1}{4}\right)}{4\sqrt{2\zeta}\sqrt[4]{G}}$ is introduced to ensure that $M_{\rm eff,II}^{(2)}$ approaches the correct classical limit, specifically to satisfy
$M_{\rm eff,II}^{(2)}=M-\frac{Q^{2}}{2x}+\frac{\Lambda x^{3}}{6G}+O\left(\sqrt{\zeta}\right).$ (150)
Now we calculate $\Psi_{\rm eff,II}$ from Eq. (138), and obtain
$\Psi_{\rm eff,II}^{(2)}=\frac{Q}{x}+\frac{\zeta^{2}Q\left(9GQ^{2}-30GMx+20x^{2}-20\Lambda x^{4}\right)}{60x^{5}}+O\left(\zeta^{5/2}\right).$ (151)
According to this result, the quantum gravity effects modify the appearance of the black hole charge in Coulomb’s potential.
III.4.3 Case of $\mathcal{X}(\mu_{\rm II})=\mu_{\rm II}$ and $\mathcal{Y}(\mu_{\rm II})=1$
In the case
$\mathcal{X}(\mu_{\rm II})=\mu_{\rm II},\quad\mathcal{Y}(\mu_{\rm II})=1,$ (152)
we have
$\mathcal{X}_{M}=-\frac{2\zeta^{2}G}{x^{3}},\quad\mathcal{Y}_{M}=0.$ (153)
Then, the factor $g_{\rm eff,II}$ is
$g_{\rm eff,II}^{(3)}(x)=\zeta^{2}\Lambda\ln(x).$ (154)
The solution to Eq. (108) reads
$M_{\rm eff,II}^{(3)}=$ $Mx^{-\zeta^{2}\Lambda}+\frac{Q^{2}}{2x\left(\zeta^{2}\Lambda-1\right)}+\frac{\zeta^{2}\Lambda x}{2\left(\zeta^{2}G\Lambda+G\right)}$
$+\frac{\Lambda x^{3}}{2\left(\zeta^{2}G\Lambda+3G\right)}.$ (155)
Hence, Eq. (138) leads to
$\Psi_{\rm eff,II}^{(3)}=\frac{Q}{x}+\frac{\zeta^{2}Q\Lambda}{x}[1+\ln(x)]+O(\zeta^{3}).$ (156)
III.5 Solutions to effective dynamics from $M_{\rm eff,III}$
The third solution $M_{\rm eff,III}$ to Eq. (85) is given by [42]
$M_{\rm eff,III}=$ $\frac{(s_{1})^{3/2}}{2G\zeta^{2}}\sin\left[\frac{\zeta^{2}}{s_{1}}\left(1+(s_{2})^{2}-\frac{(s_{4})^{2}}{4}\right)\right]$ (157)
$=$ $\frac{(E^{1})^{3/2}}{2G\zeta^{2}}\sin\left[\frac{\zeta^{2}}{E^{1}}\left(1+(K_{2})^{2}-\frac{(\partial_{x}E^{1})^{2}}{4(E^{2})^{2}}\right)\right].$
The corresponding $\mu$ is
$\mu_{\rm III}$ $=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{(s_{1})^{3}}$
$=1-\frac{4\zeta^{4}(GM_{\rm eff,III})^{2}}{x^{6}}.$ (158)
In the Schwarzschild gauge (67), Eq. (107a) yields
$x\cos\left[\frac{\zeta^{2}}{x^{2}}\left(1-\frac{x^{2}}{(E^{2})^{2}}+(K_{2})^{2}\right)\right]K_{2}=0.$ (159)
Combing Eqs. (159) and (157), we obtain
$E^{2}$ =±$=\pm\frac{x}{\sqrt{1-\frac{x^{2}}{\zeta^{2}}\left[n\pi+(-1)^{n}\arcsin\left(\frac{2\zeta^{2}GM_{\rm eff,III}}{x^{3}}\right)\right]}}$
±$\equiv\pm\frac{x}{\sqrt{f_{\rm eff,III}(x)}}.$ (160)
Using Eq. (107b), we have
$N^{2}$ $=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]f_{\rm eff,III}(x)$
$\equiv e^{g_{\rm eff,III}(x)}f_{\rm eff,III}(x).$ (161)
Then, the effective line element (81) reads
${\rm d}s^{2}$ $=-e^{g_{\rm eff,III}(x)}f_{\rm eff,III}(x){\rm d}t^{2}+\frac{{\rm d}x^{2}}{\mu_{\rm III}(x)f_{\rm eff,III}(x)}+x^{2}{\rm d}\Omega^{2}.$ (162)
Similarly, in the PG gauge (73), combing Eqs. (107a) and (157), we obtain
$N^{x}$ =±$=\pm N\sqrt{\mu_{\rm III}(x)\left[1-f_{\rm eff,III}(x)\right]}.$ (163)
Using Eq. (107b), we have
$N^{2}$ $=\exp\left[-\int\frac{x^{4}\Lambda\mathcal{X}_{M}(x)+GQ^{2}\mathcal{Y}_{M}(x)}{Gx^{2}}{\rm d}x\right]$
$=e^{g_{\rm eff,III}(x)}.$ (164)
Thus, the effective line element (81) reduces to
${\rm d}s^{2}=$ $-e^{g_{\rm eff,III}(x)}{\rm d}\tau^{2}$
$+\frac{1}{\mu_{\rm III}(x)}\left({\rm d}x\pm\sqrt{e^{g_{\rm eff,III}(x)}\mu_{\rm III}(x)\left[1-f_{\rm eff,III}(x)\right]}{\rm d}\tau\right)^{2}$
$+x^{2}{\rm d}\Omega^{2}.$ (165)
The remaining calculations and the results are the same as those in Sec. III.4, just by replacing subscript II by III.
IV Covariance of gravity coupled to a general matter field
In this section, let us discuss the covariance of gravity coupled to a general matter field.
As did for the EM field, to maintain covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$, the total Hamiltonian constraint needs to share the similar algebraic structure as Eq. (93), and to satisfy the covariance conditions presented in Sec. III.1. Let us assume that this can be implemented by modifying the classical matter Hamiltonian into the effective one $H^{\rm matter}_{\rm eff}$, which incorporates the quantity $\mu$. Since the classical matter Hamiltonian usually depends only on $E^{1}$ and $E^{2}$ as the components of the spatial metric, incorporating $\mu$ into $H^{\rm matter}_{\rm eff}$ does not introduce a dependence on derivatives of $K_{1}$. Thus, the covariance condition (i) is satisfied. Moreover, it can be verified that
$\left\{\mu,\alpha Nf(\mu,E^{1},E^{2})\right\}=\alpha\left\{\mu,Nf(\mu,E^{1},E^{2})\right\},$ (166)
for an arbitrary function $f$. As a consequence, we get
$\left\{\mu,H_{\rm eff}^{\rm matter}[\alpha N]\right\}=\alpha\left\{\mu,H_{\rm eff}^{\rm matter}[N]\right\}.$ (167)
Since Eq. (167) remains valid with replacing $H_{\rm eff}^{\rm matter}$ by $H_{\rm eff}^{GR}$ given in Eq. (92). It is concluded that Eq. (167) is satisfied by the total effective Hamiltonian constraint $H_{\rm eff}^{GR}+H_{\rm eff}^{\rm matter}$. Namely, the covariance condition (ii) is also satisfied.
According to the above discussion, to maintain the covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$ so that the total Hamiltonian constraint retains the similar algebraic structure as Eq. (93).
It is important to note that in the coupling model with matter fields, maintaining covariance also requires the matter fields themselves to be covariant. This needs us to define an effective 4-D matter field that may depend on phase space variables and the Lagrange multiplier, similar to the treatment of the field $A^{(\mu)}_{\rho}$. In general, the precise definition of the effective matter field requires further investigation.
V Summary and discussion
In this paper, we successfully extended the approach introduced in [41, 42] for restoring covariance in the effective quantum BH models to the electro-vacuum case with a cosmological constant. In Sec. II, we first recalled the Hamiltonian framework of gravity coupled to EM field in the spherically symmetric case. Specifically, we introduced the EM diffeomorphism constraint (9) following Refs. [46, 47], such that the total diffeomorphism constraint generates the spatial diffeomorphism transformations for gravity and EM field. Additionally, the EM Gauss constraint is incorporated into the Hamiltonian (7) of EM field, ensuring that the total Hamiltonian and diffeomorphism constraints of the coupling model form a subalgebra (11d)–(11f) as that of the vacuum model in [41, 42]. As a result, the analysis for spacetime covariance in the coupling model will be parallel to that in vacuum case. The covariance condition (26) for gravity and the one (27) for EM field have been examined in the classical theory. By solving the Gauss constraint, we derived the solutions (72) and (76) to the EOM in different gauge fixings, corresponding to the RN-(anti-)de Sitter metric.
In Sec. III, we studied covariance of the effective coupling model. To ensure validity of the constrain algebra (80) and the covariance conditions (i) and (ii), the classical Hamiltonian (7) of EM field has been modified into the effective one (92) by incorporating $\mu$, leading to the covariance of effective metric. To require the effective vector potential satisfying the covariance equation (96), we obtained an explicitly form (99) of $A_{\rho}^{(\mu)}$ by a detailed derivation. Based on the covariance effective coupling model, we further explored its exact solutions. Thanks to the three solutions $M_{\rm eff,I}$, $M_{\rm eff,II}$, and $M_{\rm eff,III}$ to Eq. (85) for building the effective Hamiltonian constraints of gravity in vacuum model, the corresponding exact solutions for the effective coupling model have been founded in Sec. III.3, Sec. III.4, and Sec. III.5. In the solutions corresponding to $M_{\rm eff,I}$, the vector potential $A_{\rho}^{(\mu_{\rm I})}$ remains its classical form (66) in a specific gauge choice, while the effective metric receives the quantum-corrections with the line element (120) in Schwarzschild coordinates and (123) in PG ones. Interestingly, in the solutions corresponding to $M_{\rm eff,II}$ (and $M_{\rm eff,III}$), there exist effective vector potentials that incorporate quantum gravity effects, as shown in Eqs. (151) and (156), due to existence of arbitrary functions $\mathcal{X}(\mu)$ and $\mathcal{Y}(\mu)$. Hence, in these cases, quantum gravity effects could affect EM field, besides changing the structure of spacetime.
Finally, how to extend covariance to a general matter field coupling was discussed. The main result can be summarized as follows: to preserve covariance for the effective metric $g_{\rho\sigma}^{(\mu)}$ in the coupling model with a general matter field whose classical Hamiltonian depends only on $E^{1}$ and $E^{2}$, it is sufficient to modify the classical matter Hamiltonian by incorporating $\mu$, ensuring that the total Hamiltonian constraint maintains the similar algebraic structure as Eq. (93).
Let us conclude this section by comparing our work with the study presented in [51]. The work [51] considered the scalar constraint and the vector constraint as the basis of the constraint algebra. The scalar constraint is a linear combination of the Hamiltonian constraint and the Gauss constraint considered in this paper, while the vector constraint corresponds to a linear combination of the diffeomorphism constraint and the Gauss constraint. The scalar and vector constraints form a subalgebra with the same structure as given in Eqs. (11d)–(11f). However, the vector constraint does not generate diffeomorphism transformation for $A_{\rho}$, but generates the diffeomorphism transformation for the field strength tensor $F_{\mu\nu}$. This is actually easy to explain. The vector constraint generates a diffeomorphism transformation plus a $U(1)$ gauge transformation while $F_{\mu\nu}$ is $U(1)$-gauge invariant. Due to the above facts, $F_{\mu\nu}$ is taken as a basic variable in the work [51] when constructing the effective model. Nevertheless, as demonstrated in this work, this substitution is not necessary. When the diffeomorphism constraint is directly incorporated, it is sufficient to modify the scalar constraint into the Hamiltonian constraint while ensuring that the algebraic structure given in Eqs. (11d)–(11f) is preserved. Since the vector constraint generates the diffeomorphism transformation for $A_{\rho}$, it can be chosen as a fundamental variable in the effective model. Our approach places the classical and effective models on an equal footing. Moreover, it can be easily generalized to non-Abelian Yang-Mills theory, where $F_{\mu\nu}$ is no longer gauge invariant.
Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 12275022.